OpenStudy (anonymous):

decimal as an improper fraction. 2.18

OpenStudy (anonymous):

what is the fraction for .18?

OpenStudy (anonymous):

and what is the fraction for 2

OpenStudy (anonymous):

18/10 and 2/1

OpenStudy (anonymous):

they need to be the same denominator to make this easier.

OpenStudy (anonymous):

.18 is 18/100

OpenStudy (anonymous):

so what fraction over 100 gives you 2?

OpenStudy (anonymous):

add the two together and reduce.

OpenStudy (anonymous):

so 18/100 + 200/100 = 218/100 = 109/50 good luck.